Evaluate each equation. Give exact values without a calculator. 1.cos 405

Answer by lwsshak3(11628)

Evaluate each equation. Give exact values without a calculator. 1.cos 405
**
Identity: cos(s+t)=cos s cos t-sin s sin t
cos(405)=cos(135+270)=cos 135*cos 270-sin 135*sin 270
cos 135=-√2/2
cos 270=0
sin 135=√2/2
sin 270=-1
cos 135*cos 270-sin 135*sin 270
=-√2/2*0-√2/2*(-1)
=0+√2/2
=√2/2
cos(405)=√2/2